OpenStudy (anonymous):

need help right now! please help me. Graph the two equations to solve the system. What is the solution? 1/2 x - 2y = 4 and 4y - x = -8 (-2, 0) no solution (3, 4) infinitely many solutions

OpenStudy (anonymous):

(0,-2)

OpenStudy (anonymous):

thats what i thought to but its not the correct anwser

OpenStudy (anonymous):

oh, shoot, sorry, it's wrong....in a way

OpenStudy (anonymous):

itsOk.

OpenStudy (anonymous):

they're the same line so of course plugging it in would work, but all solutions work

OpenStudy (anonymous):

infinitely many

OpenStudy (anonymous):

when rearranged it they both become y=1/4x-2

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

welcome!

OpenStudy (anonymous):

do you know this one? if not im just guna guess lol Graph the two equations to solve the system. What is the solution? y = 2x - 4 and 2x + 3y = -12 (0, -4) (4, 4) no solution infinitely many solutions

OpenStudy (anonymous):

i know its not infint solutions

OpenStudy (anonymous):

the first choice! definitely not infinite :P lol

OpenStudy (anonymous):

ok Great! :)